OpenStudy (anonymous):

find the x coordinate of the point where the graph of the function f(x)=x(sqrt(1-3x)) has a horizontal tanget line

OpenStudy (anonymous):

It has a horizontal tangent line when the point in the curve is a stationary point. So go ahead to differentiate f(x) f(x)=x(sqrt(1-3x)) \[\frac{df(x)}{dx}= \frac{2-9x}{2\sqrt{1-3x}}\] Let df(x)/dx = 0 to find the value of x that is equals to the stationary point. 0= (2-9x)/ (2srt(1-3x)) [since denominator cannot be equals to 0] 2-9x= 0 x= 2/9

OpenStudy (anonymous):

where did you get the numerator?

OpenStudy (anonymous):

It is because of the differentiation.

OpenStudy (anonymous):

how does that differentiate to get what you got for df(x)/dx=???

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+x%28sqrt%281-3x%29%29 Go the part at differentiation and click show steps.